OpenStudy (anonymous):

"Use synthetic substitution to find f (-2) and f (3)" 1) f (x) = x^2 - 2x + 6 2) f(x) = x^3 - 6x - 5 please please help me & I will return the favor. This is for my Math 3 class, we are doing Polynomials

OpenStudy (anonymous):

it is almost impossible to write synthetic division here although i have seen some attempt at it you don't need it however, since \[f(-2)=(-2)^2-2\times(-2)+6\] and \[f(3)=3^2-2\times 3+6\] and both of those are easy to compute

OpenStudy (anonymous):

Thank you for much for replying! Once I substitute the given numbers into the equation, can this be solved with a calculator?

OpenStudy (anonymous):

yes although pencil and paper will work as well none of this arithmetic is very hard

OpenStudy (anonymous):

as a matter of fact you can do much of it in your head \[(-2)^2-2\times (-5)+5=4+4+5=...\]not much to that one

OpenStudy (anonymous):

Thank you very much! Your answers were very helpful. I was over-thinking the process, I thought it was more complicated.